# Non-autoregressive Generative Models for Reranking Recommendation

### v1

@@ 3.3. Non-autoregressive sequence generation
To address the challenges mentioned earlier, non-autoregressive sequence generation takes a new approach by removing autoregressive dependencies from existing models Each element’s distribution $p(y_{t})$ depends solely on the candidates $X$:
(5) $p_{\text{NAR}}(Y|X;\theta)=\prod_{t=1}^{m}p(y_{t}|x_{1:n};\theta).$
@@
(6) $\mathcal{L}_{\text{NAR}}=-\log p_{\text{NAR}}(Y|X;\theta)=-\prod_{t=1}^{m}\log p(y_{t}|x_{1:n};\theta).$
Despite the removal of the autoregressive structure, the models retain an explicit likelihood function. The training of models employs separate cross-entropy losses for each output distribution. Crucially, these distributions can be computed concurrently during inference, introducing a marked departure from the sequential nature of autoregressive decoding. This non-autoregressive approach mitigates inference latency, thereby enhancing the efficiency of recommendation systems in real-world applications.
@@ 4. Approach
We will first introduce our non-autoregressive model structure in detail, and then detail the techniques we proposed to solve the challenges to apply it in reranking, including unlikelihood training, contrastive decoding, and sequence evaluation.
@@ 4.1. Matching model
Non-autoregressive models face challenges in converging during training due to two main reasons. Firstly, the sparse nature of training sequences makes learning difficult. Unlike text sequences that often share linguistic structures, recommended lists in training samples rarely have the same exposures, making effective learning from limited data challenging. Secondly, in the reranking stage, the same index for candidates can represent different items, resulting in a variable vocabulary as the candidates to be ranked vary across samples. Conventional models may struggle to handle such variations efficiently. To tackle these challenges, we introduce two key components to our models: a candidates encoder for effectively encoding representations of candidates and a position encoder to capture position-specific information within the target sequence.
@@
To start, we randomly initialize an embedding for each position in the target sequences. Notably, we share these position embeddings across training data to enhance learning on sparse data. We then incorporate bidirectional self-attention and cross-attention modules to learn representations for each position, leveraging information from the candidates.
@@
Additionally, to address the challenge posed by the dynamic vocabulary arising from variations in candidates across samples, we employ a matching mechanism. Specifically, we match the representation of each candidate with each position in the target sequence, resulting in the computation of probabilities for each candidate at every position.
Taking a user and the corresponding list of candidate items as input, our designed model aims to output a probability matrix $\hat{P}\in\mathbb{R}^{n\times m}$, where $\hat{P}_{i,j}$ indicates the matching score of placing the $i$-th item in the $j$-th position of the final list. Specifically, the model consists of a profile embedding layer, a candidates encoder, and a position encoder. The input profiles are first sent to the profile embedding layer and embedded into the initial candidates embeddings. Then, the candidates encoder further processes these initial embeddings via a self-attention mechanism to capture the mutual influence between items. Alongside this, the position encoder also processes a set of learnable position vectors in the same way. Finally, the output embeddings of the candidates encoder and the position encoder are used to compute the score matrix by matrix multiplication.
In the following, we describe the structure design of each model component in detail.
Embedding layer. Following previous works, we parameterize the available profiles into vector representations for users and items. Given a user $u$, associated with sparse features $x_{u}^{s}$ and dense features $x_{u}^{d}$, we embed each sparse feature value into $d$-dimensional space, while handling dense features standardization or batch normalization to ensure normal distribution. Subsequently, each user $u$ can be represented as $x_{u}\in\mathbb{R}^{|x_{u}^{s}|\times d+|x_{u}^{d}|}$, where $|x_{u}^{s}|$ and $|x_{u}^{d}|$ denote the size of sparse and dense feature space of user $u$ respectively. Similarly, we represent each item $i$ as $x_{i}\in\mathbb{R}^{|X_{i}^{s}|\times d+|x_{i}^{d}|}$. Then the representation for item candidates is denoted as $X=\{x_{1},x_{2},\cdots,x_{n}\}$
Also, we randomly initialize an embedding for each position $j$ as $p_{j}\in\mathbb{R}^{d}$. Note that these embeddings are learnable during the training process, and are fixed when using the model to evaluate the quality of the candidate list. The representation for positions is denoted as $P\in\mathbb{R}^{m\times d}$.
@@
Candidates Encoder. The candidates encoder adopts the standard Transformer architecture by stacking $N$ Transformer layers. In each layer, the architecture mainly consists of a self-attention block and a feed forward network. Denoted the input feature as $\mathbf{X}$, the self-attention block first transform it into query ($\mathbf{Q}$), key ($\mathbf{K}$) and value ($\mathbf{V}$) matrices via linear projection, as follows:
(7) $\mathbf{Q}=\mathbf{X}\mathbf{W}^{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}^{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}^{\mathbf{V}},$
@@
(8) $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})={\rm Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}})\mathbf{V},$
where $d_{k}$ denotes the dimension of the $\mathbf{K}$ matrix. Finally, it applies a linear projection to obtain the output of the self-attention block.
The feed forward network is typically placed after the self-attention block. It consists of two linear transformation layers and a non-linear activation function. It receives the output features $\mathbf{X}$ from the self-attention block, and processes them as:
(9) $FFN(\mathbf{X})=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{X}),$
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ denote the weight matrices of the two linear layers.
Position Encoder. The position encoder adopts a similar Transformer architecture as the candidates encoder. The only difference between them is that the position encoder inserts a cross attention block between the self-attention block and the feed forward network in each Transformer layer. As can be seen in Fig. 2, in each layer, the cross attention block receives the output features from the self-attention blocks in the same layer of both encoders and processes them via cross attention operation. Specifically, denoted $\mathbf{X}^{\mathbf{CE}}$ as the feature from the candidates encoder, and $\mathbf{X}^{\mathbf{PE}}$ as the feature from the position encoder, similar to the self-attention block, it first applies the linear projection to them:
(10) $\mathbf{Q}=\mathbf{X}^{\mathbf{CE}}\mathbf{W}^{\mathbf{Q}},\mathbf{K}=\mathbf{X}^{\mathbf{PE}}\mathbf{W}^{\mathbf{K}},\mathbf{V}=\mathbf{X}^{\mathbf{PE}}\mathbf{W}^{\mathbf{V}}.$
Then, it applies the formula in Eq. 8 to $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ to obtain the output feature. The cross attention uses the features from the candidates encoder as query and features from the position encoder as key and value, which can establish a connection between the embeddings of candidate items and the position.
Probability Matrix. Given the output features of the candidates encoder (denoted as $\{\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{n}}\}$) and position encoder (denoted as $\{\mathbf{p_{1}},\mathbf{p_{2}},...,\mathbf{p_{m}}\}$), we apply a matrix multiplication to them to compute the matching scores. A column-wise softmax function is then adopted to further normalize the scores. Formally, the matching score of placing the $i$-th candidate item to the $j$-th position is computed as:
(11) $\hat{p_{ij}}=\frac{exp(\mathbf{x_{i}}^{\intercal}\mathbf{p_{j}})}{\sum_{i=1}^{n}exp(\mathbf{x_{i}}^{\intercal}\mathbf{p_{j}})}$
@@
Training objectives Our model is optimized via cross-entropy loss function, which is defined as follows:
(12) $\mathcal{L}(Y,X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(\hat{p_{ij}})$
Figure 2. Architectural Overview of the Generator and Evaluator Models. The evaluator evaluates multiple sequences generated by the generator and estimates listwise scores to select the optimal sequence.
@@ 4.2. Unlikelihood Training
The divergence between text and item sequences hinders the direct application of generative models from natural language processing for item sequence recommendation. This distinction originates from the distinctive features of user interactions in recommendation scenarios. The varied feedback within recommendation sequences differs from the structured nature of natural language text. While text sequences typically adhere to standard language structures, conveying information or forming a coherent narrative, user feedback in recommendation sequences is diverse, encompassing both positive and negative interactions. Users manifest preferences through actions like clicks or likes, contributing to a diverse and nuanced feedback landscape.
Hence, the difference in objectives between maximum likelihood training(as in eq. 5) and reranking (as in eq. 2) poses a significant challenge. Although maximum likelihood training effectively captures patterns in text sequences, its applicability diminishes in recommendation scenarios where user preferences are dynamic and subjective. The essence of high-quality recommendations lies not just in sequence patterns from training data but, more crucially, in the user utility of the recommended list. User interactions with recommended items are subjective and context-dependent, introducing complexity in aligning the training goal with the desired model behavior. To navigate this challenge, we propose unlikelihood training to guide the model to assign lower probabilities to undesirable generations. This adjustment aligns the training process with the intricate feedback patterns inherent in recommendation scenarios.
Unlikelihood training decreases the model’s probability of generating a negative sequence. Given candidates $X$ and a negative sequence $Y$, the unlikelihood loss is:
(13) $\mathcal{L}_{\text{ul}}(Y_{neg},X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij})$
@@
Unlike regular text generation, where messages are clear and content-focused, the manageability of attributes like topic, style, and sentiment in the output text is evident and systematically controlled(Li et al., 2020; Welleck et al., 2019). However, recommendation sequences involve user feedback with implicit signals. For instance, a lack of interaction with a recommended item may suggest disinterest. This highlights the model’s need to understand both explicit and implicit cues in user feedback. Effective control over a generation in recommendation sequences becomes crucial to tailor the output based on users’ preferences and behaviors, thus enhancing the personalized nature of recommendations. Specifically, we regard a sequence as a positive or negative sequence based on the overall utility defined in section 3.1, and the corresponding loss is as follows:
(14) $\mathcal{L}_{\text{ul}}(Y,X)=\left\{\begin{array}[]{ll}-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij})&\text{ if }\mathcal{R}(u,Y)<\alpha\\ -\sum_{i=1}^{n}\sum_{j=1}^{m}P_{ij}\log(\hat{P_{ij}})&\text{ if }\mathcal{R}(u,Y)\geq\alpha\end{array}\right.$
@@ 4.3. Contrastive Decoding
Compared with autoregressive generation, the non-autoregressive approach significantly enhances computation efficiency and makes it feasible to deploy in real-time recommendation systems. However, non-autoregressive generation uses the conditional independence assumption: each target item’s distribution $p(y_{t})$ depends only on the candidates $X$. This deviation from autoregressive models challenges in capturing the inherently multimodal nature of the distribution of valid target sequences. Take machine translation for example, when translating the phrase ”thank you” into German could result in multiple valid translations such as ”Vielen Dank” and ”Danke”. However, non-autoregressive models may generate unplausible translations like ”Danke Dank”. The conditional independence assumption in eq. 5 restricts the model’s ability to effectively capture the multimodal distribution in the set of possible target sequences.
@@
Essentially, the assumption of conditional independence limits the model’s ability to navigate a vast solution space and identify the most suitable permutation from numerous valid options for a given set of candidates. This limitation is especially notable in scenarios where the number of reasonable target sequences far exceeds those encountered in natural language processing. Consequently, non-autoregressive frameworks face the challenge of addressing the impact of conditional independence to improve their capacity for generating diverse and contextually appropriate target sequences. To tackle this, we propose contrastive decoding to model the co-occurrence relationship between items and thereby improve the target dependency.
@@
Contrastive decoding incorporates a diversity prior that regulates the sequence decoding procedure with the predicted distribution of the model. This is based on the intuition that a good recommended list needs to be composed of a wide variety of items. In fact, contrastive decoding leverages a similarity score function as a regularizer when decoding, which captures the dependence between different positions in the target sequence.
@@
Formally, given the previous context $y_{<t}$, at time step $t$, the selection of the output $y_{t}$ follows
(15) $y_{t}=argmax_{x\in X}(1-\alpha)\times p(x|p_{i},X)-\alpha\times max(c(\mathbf{x}_{i},\mathbf{x}_{j}))$
@@
where $11\leq j\leq t-$ 1. In eq. 15, the first term, model confidence is the probability of candidates $x$ predicted by the model. The second term, similarity penalty, measures how discriminative of candidate $x$ with respect to the previously selected items, where $c(\mathbf{x_{i}},\mathbf{x_{j}})$
(16) $c(\mathbf{x_{i}},\mathbf{x_{j}})=\frac{\mathbf{x_{i}}^{\top}\mathbf{x_{j}}}{\|\mathbf{x_{i}}\|\cdot\|\mathbf{x_{j}}\|}.$
Specifically, the similarity penalty is defined as the maximum similarity between the representation of $x$ and all items in $y_{<t}$. Intuitively, a larger similarity penalty it is more similar to the context.
@@
Also, to encourage the language model to learn discriminative and isotropic item representations, we incorporate a contrastive objective into the training of the language model. Specifically, given a sequence $X$, the $\mathcal{L}_{\textup{CL}}$ is defined as
(17) $\mathcal{L}_{\text{position}}=\frac{1}{n\times(n-1)}\sum_{i=1}^{n}\sum_{j=1,j\neq i}^{n}\max\{0,\rho-c(\mathbf{x}_{i},\mathbf{x}_{i})+c(\mathbf{x}_{i},\mathbf{x}_{j})\},$
@@
where $\rho\in[-1,1]$ is a pre-defined margin and $x_{i}$ is the representation of item $x_{i}$ produced by the model.
(18) $\mathcal{L}_{\text{item}}=\frac{1}{m\times(m-1)}\sum_{i=1}^{m}\sum_{j=1,j\neq i}^{|\boldsymbol{x}|}\max\{0,\rho-c(\mathbf{p}_{i},\mathbf{p}_{i})+c(\mathbf{p}_{i},\mathbf{p}_{j})\},$
@@
Intuitively, by training with $\mathcal{L}_{\textup{CL}}$, the model learns to pull away the distances between representations of distinct tokens. [^1] By definition, the cosine similarity $s(h_{x_{i}},h_{x_{i}})$ of the identical token $x_{i}$ is $1.0$. Therefore, a discriminative and isotropic model representation space can be obtained. The overall training objective $\mathcal{L}$ is then defined as
@@
where the maximum likelihood estimation (MLE) objective $\mathcal{L}$ is described in eq. 6. Note that, when the margin $\rho$ in $\mathcal{L}_{\textup{CL}}$ equals to 00, the $\mathcal{L}_{\textup{SimCTG}}$ degenerates to the vanilla MLE objective $\mathcal{L}_{\textup{MLE}}$.

### v3

@@ 3.3. Non-autoregressive sequence generation
To address the aforementioned challenges, non-autoregressive sequence generation eliminates autoregressive dependencies from existing models. Each element’s distribution $p(y_{i})$ depends solely on the candidates $X$:
(5) $p_{\text{NAR}}(Y|X;\theta)=\prod_{i=1}^{m}p(y_{i}|x_{1:n};\theta).$
@@
(6) $\mathcal{L}_{\text{NAR}}=-\log p_{\text{NAR}}(Y|X;\theta)=-\prod_{i=1}^{m}\log p(y_{i}|x_{1:n};\theta).$
Despite the removal of the autoregressive structure, the models retain an explicit likelihood function. The training of models employs separate cross-entropy losses for each output distribution. Crucially, these distributions can be computed simultaneously during inference, which significantly differs from the sequential process of autoregressive models. This non-autoregressive approach reduces inference latency, thereby enhancing the efficiency of recommendation systems in real-world applications.
@@ 4. Approach
In this section, we present a detailed introduction of NAR4Rec. We will first discuss our non-autoregressive model structure, which estimates the probability by a matching model in section 4.1. Then, we delve into unlikelihood training, a method aimed at discerning feedback within the recommended sequence in section 4.2. Finally, we propose contrastive decoding to model the dependency in target sequence in section 4.3. The sequence evaluator in our generator-evaluator framework is introduced in section 4.4.
@@ 4.1. Matching model
Figure 2. Architectural Overview of the Generator and Evaluator Models. The evaluator evaluates multiple sequences generated by the generator and estimates listwise score to select the optimal sequence.
@@
Non-autoregressive models encounter challenges in training convergence due to two main reasons. Firstly, the sparse nature of training sequences presents learning difficulties. Unlike text sequences that often share linguistic structures, recommended lists in training samples seldom have the same exposures, posing challenges for effective learning from limited data. Secondly, during the reranking stage, identical index for candidates may denote different items, leading to a variable vocabulary as the candidates to be ranked vary across samples.
@@
Conventional models may struggle to handle such variations efficiently. To tackle these challenges, we introduce two key components to our models: a candidates encoder for effectively encoding representations of candidates and a position encoder to capture position-specific information within the target sequence. Initially, we randomly initialize an embedding for each position in the target sequences. Notably, we share these position embeddings across training data to enhance learning on sparse data. Subsequently, we integrate bidirectional self-attention and cross-attention modules to acquire representations for each position, leveraging information from the candidates.
@@
Additionally, to address the challenge posed by the dynamic vocabulary arising from variations in candidates across samples, we employ a matching mechanism. Specifically, we match each candidate with each position in the target sequence, thereby yielding probabilities for each candidate at every position. In the following, we elaborate on the structure of NAR4Rec.
Given a user $u$ and candidates $X=\left\{x_{1},x_{2},\cdots,x_{n}\right\}$, the hidden representation of $x_{i}$ is $\mathbf{x}_{i}\in\mathbb{R}^{d_{x}}$. We stack $\mathbf{x}_{i}$ together into matrix $\mathbf{X}\in\mathbb{R}^{n\times d_{x}}$. Additionally, we randomly initialize an embedding vector for each position $j$ as $t_{j}\in d_{t}$. Also, we stack $t_{j}$ into $\mathbf{T}\in\mathbb{R}^{m\times d_{t}}$. To align the embedding dimension of $\mathbf{X}$ with that of $\mathbf{T}$, we project them into same hidden dimension $d$ by a linear projection layer. Then $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{T}\in\mathbb{R}^{m\times d}$. Consequently, $\mathbf{X}$ is represented as $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{T}$ as $\mathbf{T}\in\mathbb{R}^{m\times d}$.
@@
Candidates Encoder. The candidates encoder adopts the standard Transformer architecture(Vaswani et al., 2017) by stacking $L$ Transformer layers. In each layer, the architecture mainly consists of two blocks, a self-attention block and a feed-forward network. An input $\mathbf{X}$ for self-attention block is linearly transformed into query ($\mathbf{Q}_{x}$), key ($\mathbf{K}_{x}$) and value ($\mathbf{V}_{x}$) as follows:
(7) $\mathbf{Q}_{x}=\mathbf{X}\mathbf{W}^{\mathbf{Q}}_{x},\mathbf{K}_{x}=\mathbf{X}\mathbf{W}^{\mathbf{K}}_{x},\mathbf{V}_{x}=\mathbf{X}\mathbf{W}^{\mathbf{V}}_{x},$
@@
(8) $\text{Attention}(\mathbf{Q}_{x},\mathbf{K}_{x},\mathbf{V}_{x})={\rm Softmax}(\frac{\mathbf{Q}_{k}\mathbf{K}_{x}^{T}}{\sqrt{d}})\mathbf{V}_{x},$
For a multi-head version of self-attention mechanism, the input is linearly projected into $\mathbf{Q}_{x}$, $\mathbf{K}_{x}$ and $\mathbf{V}_{x}$ with $h$ times using individual linear projections to small dimensions(e.g. $d_{k}=\frac{d}{h}$). Finally, the output of self-attention(SAN) is
(9) SAN $=[head_{1},\cdots,head_{h}]\mathbf{W}^{O}_{x},$
$head_{i}$ $=Attention(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}).$
The feed-forward network is typically placed after the self-attention block,
(10) $FFN(\mathbf{X})=\sigma(\mathbf{X}\mathbf{W}^{in}_{x})\mathbf{W}^{out}_{x},$
where $\mathbf{W}^{out}_{x}$ and $\mathbf{W}^{in}_{x}$ denote the weight matrices of the two linear projection.
Position Encoder. The position encoder adopts a similar Transformer architecture as the candidates encoder. The key difference between them is that the position encoder inserts a cross-attention block between the self-attention block and the feed-forward network in each Transformer layer. As can be seen in Fig. 2, in each layer, the cross-attention block receives the hidden representation from the self-attention blocks of both encoders and processes them via cross-attention operation. Specifically, the hidden representation from the candidates encoder and position encoder are denoted as $\mathbf{X}$ and $\mathbf{T}$, respectively. Similar to the self-attention block, we initially apply linear projection to them:
(11) $\mathbf{Q}=\mathbf{T}\mathbf{W}^{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}^{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}^{\mathbf{V}}.$
Then, we applies the formula in eq. 9 to $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ to get the output hidden representation. The cross-attention is introduced to capture the correlation between candidates and target sequence.
Probability Matrix. To compute the probability matrix, we perform a matrix multiplication on the output hidden representation from the candidates encoder (denoted as $\{\mathbf{x_{1}},\mathbf{x_{2}},...,\mathbf{x_{n}}\}$) and position encoder (denoted as $\{\mathbf{t_{1}},\mathbf{t_{2}},...,\mathbf{t_{m}}\}$). Subsequently, we apply a column-wise softmax function to normalize the scores. Formally, the probability score of placing the $i$-th candidate item to the $j$-th position is calculated as:
(12) $\hat{p_{ij}}=\frac{exp(\mathbf{x_{i}}^{\intercal}\mathbf{t_{j}})}{\sum_{i=1}^{n}exp(\mathbf{x_{i}}^{\intercal}\mathbf{t_{j}})}.$
Training objectives NAR4Rec is trained via cross-entropy loss function, defined as follows:
(13) $\mathcal{L}(Y,X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(\hat{p_{ij}}),$
@@ 4.2. Unlikelihood Training
The discrepancy between text and item sequences hinders the direct application of generative models from text to item recommendation. This disparity arises from the unique characteristics of user interactions in recommendation scenarios. Unlike the structured nature of natural language text, the feedback within recommendation sequences is diverse due to the varied nature of user interactions. While text sequences typically follow conventional language structures to convey information or construct a coherent narrative, user feedback in recommendation sequences is characterized by diverse actions such as clicks or likes, reflecting a diverse and nuanced feedback.
Consequently, the difference in objectives between maximum likelihood training(as in eq. 5) and reranking (as in eq. 2) poses a significant challenge. Although maximum likelihood training effectively captures patterns in text sequences, its applicability diminishes in recommendation scenarios where user preferences are dynamic and subjective. The essence of high-quality recommendations lies not just in sequence patterns from training data but, more crucially, in the user utility of the recommended list. User interactions with recommended items are subjective and context-driven, adding complexity to aligning the training objective with the desired model behavior. To address this challenge, we propose unlikelihood training, guiding the model to assign lower probabilities to undesired generations. This adjustment aligns the training process with the intricate feedback patterns.
Unlikelihood training reduces the model’s probability of generating a negative sequence. Given candidates $X$ and a negative sequence $Y_{neg}$, the unlikelihood loss is:
(14) $\mathcal{L}_{\text{ul}}(Y_{neg},X)=-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij}),$
@@
Unlike text generation, where messages are clear and content-focused, managing attributes like topic, style, and sentiment in the output text is straightforward(Li et al., 2020; Welleck et al., 2019). However, recommendation sequences involve user feedback with implicit signals. For instance, a lack of interaction with a recommended item may suggest disinterest. This highlights the model’s need to understand both explicit and implicit cues in user feedback. Effective control over generation in recommendation sequences becomes crucial to tailor the output based on user preferences and behaviors, thus enhancing the personalized recommendations. Specifically, we classify a item sequence as positive or negative based on the overall utility defined in section 3.1, and the corresponding loss is as follows:
(15) $\mathcal{L}_{\text{ul}}(Y,X)=\left\{\begin{array}[]{ll}-\sum_{i=1}^{n}\sum_{j=1}^{m}p_{ij}\log(1-\hat{p}_{ij})&\text{ if }\mathcal{R}(u,Y)<\tau\\ -\sum_{i=1}^{n}\sum_{j=1}^{m}P_{ij}\log(\hat{P_{ij}})&\text{ if }\mathcal{R}(u,Y)\geq\tau\end{array}\right.$
@@ 4.3. Contrastive Decoding
Compared with autoregressive generation, the non-autoregressive approach significantly enhances computation efficiency and makes it feasible to deploy in real-time recommendation systems. However, non-autoregressive generation introduce the conditional independence assumption: each target item’s distribution $p(y_{i})$ depends only on the candidates $X$. This deviation from autoregressive models poses challenges in capturing the inherently multimodal nature of the distribution of valid target sequences. Take machine translation for example, when translating the phrase ”thank you” into German could result in multiple valid translations such as ”Vielen Dank” and ”Danke”. However, non-autoregressive models may generate unplausible translations like ”Danke Dank”. The conditional independence assumption in eq. 5 restricts the model’s ability to effectively grasp the multimodal distribution in target sequences.
@@
Essentially, the assumption of conditional independence limits the model’s ability to navigate a vast solution space and identify the most suitable permutation from numerous valid options for a given set of candidates. This limitation is especially evident in recommendation where the number of reasonable target sequences far exceeds those encountered in text. Consequently, non-autoregressive frameworks grapple with the challenge of mitigating the impact of conditional independence to improve their capacity for generating diverse and contextually appropriate target sequences. To tackle this, we propose contrastive decoding to model the co-occurrence relationship between items and thereby improve the target dependency.
@@
Contrastive decoding incorporates a diversity prior that regulates the sequence decoding procedure. This is grounded in the intuition that an effective recommended list needs to be composed of a wide variety of items. In fact, contrastive decoding leverages a similarity score function as a regulator when decoding, capturing the interdependence between various positions in the target sequence.
@@
Formally, given the preceding context $y_{<i}$, at time step $i$, the selection of the output $y_{i}$ follows
(16) $y_{t}=argmax_{x\in X}(1-\alpha)\times p(x|p_{i},X)-\alpha\times max(s(\mathbf{x},\mathbf{x}_{j})),$
@@
where $10\leq j\leq i-$ 1. In eq. 16, the first term, termed as model confidence, denotes the probability of candidates $x$ predicted by the model. The second term, known as similarity penalty, quantifies the distinctiveness candidate $x$ concerning the previously selected items, where $s(x,\mathbf{x_{j}})$ is computed as:
(17) $s(\mathbf{t_{i}},\mathbf{t_{j}})=\frac{\mathbf{t_{i}}^{\top}\mathbf{t_{j}}}{\|\mathbf{t_{i}}\|\cdot\|\mathbf{t_{j}}\|}.$
Specifically, the similarity penalty is defined as the maximum similarity between the representation of $x$ and all items in $y_{<i}$. NAR4Rec utilizes the dot product item embedding and position embedding to compute the probability matrix. Higher embedding similarity between items often means similar probability in a certain position. We introduce such penalty to introduce intra-list correlation.
@@
Also, to encourage the language model to learn discriminative and isotropic item representations, we incorporate a contrastive objective into the training of the language model. Specifically, given a sequence $X$, the $\mathcal{L}_{\text{position}}$ and $\mathcal{L}_{\text{item}}$ are defined as:
(18) $\mathcal{L}_{\text{position}}=\frac{1}{n\times(n-1)}\sum_{i=1}^{n}\sum_{j=1,j\neq i}^{n}\max\{0,\rho-s(\mathbf{x}_{i},\mathbf{x}_{i})+s(\mathbf{x}_{i},\mathbf{x}_{j})\},$
@@
where $\rho\in[-1,1]$ is a pre-defined margin and $x_{i}$ is the hidden representation of item $x_{i}$ from candidates encoder.
(19) $\mathcal{L}_{\text{item}}=\frac{1}{m\times(m-1)}\sum_{i=1}^{m}\sum_{j=1,j\neq i}^{|\boldsymbol{x}|}\max\{0,\rho-s(\mathbf{t}_{i},\mathbf{t}_{i})+s(\mathbf{t}_{i},\mathbf{t}_{j})\},$
@@
where $t_{i}$ is the hidden representation of position $t_{i}$ from the position encoder. Intuitively, by training with $\mathcal{L}_{\textup{CL}}$, the model learns to pull away the distances between representations of distinct tokens. [^1] By definition, the cosine similarity $s(h_{x_{i}},h_{x_{i}})$ of the identical token $x_{i}$ is $1.0$. Therefore, a discriminative and isotropic model representation space can be obtained.
The overall training objective $\mathcal{L}$ is then defined as
@@
where the unlikelihood training objective is described in eq. 15. Note that, when the margin $\rho$ in $\mathcal{L}_{\text{position}}$ and $\mathcal{L}_{\text{item}}$ equals to 00, the $\mathcal{L}(Y,X)$ degenerates to the vanilla unlikelihood training objective $\mathcal{L}_{\textup{ul}}$.